# MTDT: A Multi-Task Deep Learning Digital Twin

### v1

@@ III-D Travel Time Module
Queue Length subnetwork module $M_{tt}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of travel time histogram $tt$ estimation. Here, the input data is fed in the form of Multivariate Time series $v^{\prime}_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phases, pair 7 variables including inflow waveform time series $inf$ estimated by $M_{inf}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant temporospatial characteristics of urban traffic intersections. In this setting, for simplicity reasons, only the three most effective driving behavior i.e., accel, lcCooperative, minGap in traffic dynamic of intersection as shown in [23]. The architecture of this module remains the same as the $M_{ql}$ module.
@@ III-E Loss Construction
In the training phase, the learning process for both primary and secondary tasks operates independently while being supervised by log information generated by the SUMO micro-simulator for each target variable. The input information is represented either as Simulation Graphs for GATs modules handling primary tasks or as Multivariate Time Series for CNNs modules addressing secondary tasks. While, in the inference phase, the secondary tasks are facilitated by information derived from primary tasks involved in constructing their inputs (i.e., shown by the activation of red-colored arrows in the Figure 2). The output of secondary modules in inference mode are generated by performing 1D convolutional operation on the resulting 1D Multivariate Time Series. It’s worth noting that this study employs a simplified design to illustrate the benefit of this generic approach, more complex architectures with improved feature learning might lead to better results.
To optimize the generalization of the optimization scheme within our multi-task learning paradigm, we employ appropriate normalization techniques tailored to each task to ensure equal weighting of partial loss values computed by different tasks. $M_{ext}$, $M_{inf}$, and $M_{ql}$ is optimized using the mean squared error (squared L2 norm) loss metric while $M_{tt}$ is optimized by cross-entropy loss function with the number classes equal to the number of aggregated bins in the original histogram of the distribution. The normalization technique used for $y_{ext}$ and $y_{inf}$ and $y_{tt}$ is min-max normalization, while $y_{ql}$ is log scaled. For each normalization and scaling formula, a counterpart denormalization function is utilized to transform $\{\hat{y}_{t}\}_{t\in T}$ back to the original scale before computing error values for estimations.
@@
The model is trained on the train portion, and a grid search algorithm is employed to fine-tune regularization hyperparameters. The model is evaluated on to the test partition and the error values are reported for the best-performing model across 100 epochs of training.
In this study, the secondary sub-network modules are designed with as few parameters as possible and low computational overhead. This design ensures easy integration into the primary block, enhancing the efficiency of multi-task learning within the resulting model. While we deliberately maintain a simplistic architecture to showcase the advantages of this learning paradigm in our application, introducing subtle additional complexities could potentially yield improved results through enhanced feature learning capabilities.
@@ IV Traffic Data Generation
Our dataset is based on over 400,000 simulation-hours of SUMO data. A 9-intersection urban corridor located in a large metropolitan region in the United States of America was chosen. The dataset has been described in [23] in detail. We reiterate the important points here.
@@
Building upon the configuration parameters and the range of variation detailed in [23], we augment our data generation process with additional processing. This enhancement enables the collection of several Measures of Effectiveness (MOEs) corresponding to the state of traffic flow across various simulated counterfactual traffic scenarios.
For signal timing plans, we use Ring-and-Barrier operation with a common cycle length. Parameters such as the minimum/maximum green times, yellow times, and red times for the intersections are based on signal timing sheets. However, we do vary the common cycle lengths and barrier times. A random common cycle length is chosen between 120 seconds and 240 seconds. Once this is selected, random signal cycle offsets are chosen for each intersection separately. A random barrier time is chosen. Driving behavior parameters (such as acceleration, deceleration, emergency braking, minimum gap between vehicles, headway, lane-changing parameters, etc.) are also varied. Finally with some additional processing, we get the outputs of several Measures of Effectiveness(MOEs) to measure the state of traffic flow generated at counterfactual scenarios.
Queue Length. The maximum queue length time series for each approach or traffic phase represents the instant magnitude of the maximum queue length in the unit of meter aggregated at 5-second bucket resolution. In order to get the variation of queue length over time, we take the maximum queue length seen for a group of lanes associated with each phase of movement at a 5-second resolution. This is obtained directly by parsing SUMO log files and tracking queue lengths on a per-second basis. For the through-lanes (i.e. lanes serving the even number phases 2/4/6/8), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the through-lanes and (b) 2-hop lanes w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. Broadly, this number gives the maximum possible queue length that any of the lanes belonging to the through-lanes, might experience. For the left-lanes (i.e. lanes serving the odd number phases 1/3/5/7), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the left-lanes and (b) the left-most 2-hop lane w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. The reason for choosing only the left-most 2-hop lane for (b) is that it is assumed that many vehicles that wish to take a left-turn, would have begun changing into the left-most lane in anticipation of the start of the left-turn buffer. Thus, the maximum queue length should take into account the left-most lane.
@@ V Experimental Results
In this section, we assess the performance of our proposed multi-tasking digital twin, referred to as MTDT. This MTDT model undergoes training on a dataset comprising 22,000 Simulation Records meticulously gathered from a diverse range of eight intersections. Additionally, we introduce two variants for comparison: MTDT-Single, trained on a dataset of 19,000 Simulation Records from a single intersection, and MTDT-MOE, a modified version of MTDT that excludes primary modules $M_{ext}$ and $M_{inf}$ from the pipeline, resulting in the output of only two sets of MOEs variables, namely $ql$ and $tt$, instead of the full set of variables. Lastly, we include TQAM [15] a RNN Auto-encoder as a baseline deep learning-based digital twin previously introduced in our research for the estimation of maximum queue length.
@@
Table. 2 shows our proposed digital has the potential to effectively learn and generalize intersection-level traffic simulation to any arbitrary intersection when it is exposed for an adequate number of epochs to the different intersections. MTDT-Single and MTDT show similar results for inflow and exit waveform time series and travel time distribution estimation. Although MTDT-Single outperforms in the estimation of the maximum queue length, the comparison, in general, implies that our digital twin is able to effectively generalize the joint estimation of traffic flow and MOE estimation at the intersection level regardless of its lane design and configuration (topology). The benefit of multi-task learning can be proved by comparing the performance of MTDT-MOE and MTDT-Single. The performance of secondary tasks is supported by the outputs of the primary task, resulting in better performance in MTDT-Single compared to MTDT-MOE.
The minimal marginal increase in error values for all variable remains minimal relative to the aggregation level indicates consistent reliability and robustness in the model’s performance across varying data resolutions. We also note the model exhibits enhanced performance in estimating travel time for lower quantiles of the data distribution while showing a slight improvement, particularly at the 90th quantile. This suggests that our designed digital twin is also particularly adept at capturing relatively rare or extreme patterns.
@@
TABLE 2: Performance of MTDT. The performance of our proposed digital twin is assessed across various aggregation resolutions of maximum queue length and at different quantiles of the travel time distribution. MTDT is compared to its variant (MTDT-Single) when trained on a single intersection, and MTDT-Single is further compared to MTDT-MOE, where GAT modules of primary tasks are ablated from the pipeline. This comparison is conducted in terms of NRMSE, RMSE and MAE, measured in vehicle-per-bucket for Exit and Inflow waveforms, meter-per-bucket for maximum queue length, and vehicle-per-quantile for travel time distribution. Moreover, the level of confidence for estimating the relevant variable across various resolution levels are shown as 95% confidence intervals (CI).
For the multifaceted and fine-grained results, our designed multi-tasking simulation stands as unique to its counterparts, thus not directly comparable to existing baselines. However, we attempt to contrast our findings with those obtained by TQAM designed to make an accurate estimation of queue length estimation in a specific approach of a specific topology of target intersection. To achieve this, we partition our data into the same queue length buckets as reported for this model and compare error values in terms of (percentual) Mean Absolute Error (MAE) in Table 3. The partitioning encompasses two groups for each of the low (L), medium (M), and high (H) ranges of maximum queue length. These groups, ranging from L1 to H2, encompass 28%, 34%, 19%, 11%, 5%, and 3% of data samples, respectively. Our observations indicate that both MTDT and MTDT-Single exhibit comparable performance to TQAM. Specifically, MTDT outperforms TQAM in certain partitions of data where very short (0-40 meters) or long (above 200 meters) queue length tails are recorded, while MTDT-Single consistently performs equivalently or better than TQAM across nearly all partitions.
@@
TABLE 4: Evaluation at various green time allocations. The performance of the model is evaluated at different plit of data based on corridor-through green time percentage buckets. The error values are reported based on NRMSE, MSE and MAE values measured in vehicle per 5-second bucket for Exit and Inflow waveforms, meter per 5-second bucket for maximum queue length, and vehicle per 60th percentile for travel time distribution.
@@ VI Related Work
In the realm of non-linear models, multi-kernel SVM [10] and weighted k-nearest neighbor [21] with MapReduce algorithm on Hadoop are introduced. When it comes to hybrid models, ensemble empirical mode decomposition with adaptive noise [16] and bidirectional RNN combined with road crossing vector coding [26] can be mentioned. Machine learning models in this line of research are deep learning frameworks based on the temporal convolutional network (TCN) [27], parallel computing of deep belief networks (DBNs) [25], deep characteristic learning using DBN with the multi-objective particle swarm optimization (PSO) algorithm [9]. [3] and [11] developed a deep hybrid neural network combined with CNN and LSTM using trajectory data for better accuracy and robustness of forecasting performance, while [paper[1]] combined GCN and attention mechanism for lane-wise and topology invariant traffic flow estimation.
@@ VII Conclusions and Future Work
In this paper, we demonstrate the effectiveness of multi-tasking as a learning paradigm for deep learning-based digital twins. We introduce a Multi-Task Deep Learning Digital Twin, so-called MTDT, capable of successfully performing diverse albeit related traffic simulation tasks related to an urban intersection with an arbitrary topology and characteristics. Our findings highlight the benefits of the multi-task learning approach in the design on our digital twin, not only for secondary tasks whose inputs rely on the output of primary tasks but also for primary tasks to generalize and enhance their performance through shared learned representations and joint optimization. Furthermore, we design two experiments to specifically illustrate how the model’s performance varies across different partitions of data when simulation scenarios are segmented based on signal timing plans or maximum recorded queue length. The results underscore the consistency and robustness of our approach across different experiments. Moreover, we discuss the lack of baseline due to the uniqueness and comprehensiveness of our approach. Although, we show it still outperforms in individual tasks compared to some previously introduced digital twins specialized for that specific tasks.
@@
In this approach, we use lane-wide inflow time series waveforms $inf$ estimated by the graph-based digital twin $G_{inf}$ as vehicles approach the target intersection from various directions, we construct the Origin-Destination (O-D) matrix. Subsequently, we employ the od2trips tool of SUMO to generate route files and evaluate each traffic scenario using a variety of Measures of Effectiveness (MOEs). An Optimum signal timing plan can be viewed as the most favorable trade-off among desired MOEs, considering a range of traffic conditions such as signal timing plans and driving behaviors. We consider this topic a significant avenue for our future research.
MTDT can serve leveraged as a valuable tool to assist traffic decision-makers in comprehending the potential impact of alterations to signal timing plans and proposed changes to road geometry. Additionally, it can be seamlessly integrated into signal timing optimization frameworks, including those based on reinforcement learning and other methodologies.
Moreover, this approach can be expanded from isolated intersections to corridors and networks, paving the way for traffic signal timing plan optimization, which will be a focal point of our future endeavors. Moreover, our model holds promise for integration into MapReduce frameworks, facilitating distributed training and parallel prediction of traffic flow.

### v2

@@ III-D Travel Time Module
Queue Length subnetwork module $M_{tt}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of travel time histogram $tt$ estimation. Here, the input data is fed in the form of Multivariate Time series $v^{\prime}_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phases, pair 7 variables including inflow waveform time series $inf$ estimated by $M_{inf}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant spatiotemporal characteristics of urban traffic intersections. The architecture of this module remains the same as the $M_{ql}$ module.
@@ III-E Loss Construction
In the training phase, the learning process for both primary and secondary tasks operates independently while being supervised by log information generated by the SUMO micro-simulator for each target variable. The input information is represented either as Simulation Graphs for GATs modules handling primary tasks or as Multivariate Time Series for CNNs modules addressing secondary tasks. While, in the inference phase, the secondary tasks are facilitated by information derived from primary tasks involved in constructing their inputs (i.e., shown by the activation of red-colored arrows in Figure 2). The output of secondary modules in inference mode is generated by performing a 1D convolutional operation on the resulting 1D Multivariate Time Series. It’s worth noting that this study employs a simplified design to illustrate the benefit of this generic approach, more complex architectures with improved feature learning might lead to better results.
In the multi-tasking paradigm, consistency across the scale of outputs scales is important. In this experiment, a traffic scenario for a certain intersection includes several variables: $ext$ and $inf$ are time series with a range of 0-8 in the unit of number of vehicles, while $ql$ is time series with a range of 0-1200 in the unit of meter, and $tt$ is a histogram with range 0-700 in the unit of seconds. To optimize the generalization of the optimization scheme within our multi-task learning paradigm, we employ appropriate experimental normalization techniques tailored to each task to ensure equal weighting (prioritization) of partial loss values computed by different tasks. We use the min-max normalization technique for $y_{ext}$ and $y_{inf}$ and $y_{tt}$ and log normalization for $y_{ql}$. For each normalization and scaling formula, a counterpart denormalization function is utilized to transform $\{\hat{y}_{t}\}_{t\in T}$ back to the original scale before computing error values for estimations.
@@
The choice of loss function for each task also is aligned with the data type and task objectives. We use the mean squared error loss function for time series and cross-entropy loss function for histogram optimization treating the number of bins as the number of labels for its ordinal categorical variable.
The model is trained on the train portion ($75\%$ split of data), and a grid search algorithm is employed to fine-tune regularization hyperparameters on the $15\%$ split of the validation portion. The model is evaluated on to the test partition and the error values are reported for the best-performing model across 100 epochs of training.
In this study, we deliberately maintain a simplistic architecture to showcase the advantages of this learning paradigm in our application, introducing subtle additional complexities could potentially yield improved results through enhanced feature learning capabilities.
@@ IV Traffic Data Generation
Our dataset is based on over 400,000 simulation hours of SUMO data. A 9-intersection urban corridor located in a large metropolitan region in the United States of America was chosen. We reiterate the important points here; more details on the dataset are provided in [23].
@@
Building upon the configuration parameters and the range of variation detailed in Supplementary Table 1 of [23], we augment our data generation process with additional processing. This enhancement enables the collection of several Measures of Effectiveness (MOEs) corresponding to the state of traffic flow across various simulated counterfactual traffic scenarios.
For signal timing plans, we use Ring-and-Barrier operation with a common cycle length. Parameters such as the minimum/maximum green times, yellow times, and red times for the intersections are based on signal timing sheets. However, we do vary the common cycle lengths and barrier times. A random common cycle length is chosen between 120 seconds and 240 seconds. Once this is selected, random signal cycle offsets are chosen for each intersection separately. A random barrier time is chosen. Driving behavior parameters (such as acceleration, deceleration, emergency braking, minimum gap between vehicles, headway, lane-changing parameters, etc.) are also varied. Finally, with some additional processing, we get the outputs of several Measures of Effectiveness(MOEs) to measure the state of traffic flow generated at counterfactual scenarios.
Queue Length. The maximum queue length time series for each approach or traffic phase represents the instant magnitude of the maximum queue length in the unit of meter aggregated at a 5-second bucket resolution. In order to get the variation of queue length over time, we take the maximum queue length seen for a group of lanes associated with each phase of movement at a 5-second resolution. This is obtained directly by parsing SUMO log files and tracking queue lengths on a per-second basis. For the through-lanes (i.e. lanes serving the even number phases 2/4/6/8), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the through-lanes and (b) 2-hop lanes w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. Broadly, this number gives the maximum possible queue length that any of the lanes belonging to the through-lanes, might experience. For the left-lanes (i.e. lanes serving the odd number phases 1/3/5/7), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the left-lanes and (b) the left-most 2-hop lane w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. The reason for choosing only the left-most 2-hop lane for (b) is that it is assumed that many vehicles that wish to take a left-turn, would have begun changing into the left-most lane in anticipation of the start of the left-turn buffer. Thus, the maximum queue length should take into account the left-most lane.
@@ V Experimental Results
In this section, we assess the performance of our proposed multi-tasking digital twin, referred to as MTDT. This MTDT model undergoes training on a dataset comprising 22,000 Simulation Records meticulously gathered from a diverse range of eight intersections. Additionally, we introduce two variants for comparison: MTDT-Single, trained on a dataset of 19,000 Simulation Records from a single intersection, and MTDT-MOE, a modified version of MTDT that excludes primary modules $M_{ext}$ and $M_{inf}$ from the pipeline, resulting in the output of only two sets of MOEs variables, namely $ql$ and $tt$, instead of the full set of variables. Lastly, we include TQAM [15] an RNN Auto-encoder as a baseline deep learning-based digital twin previously introduced in our research for the estimation of maximum queue length.
@@
Table. 2 shows our proposed digital twin (MTDT) has the potential to effectively learn the joint estimation of traffic flow and MOE estimation at any intersection regardless of its lane design and configuration (topology). Although MTDT-Single is trained and evaluated at a single intersection, the error values are fairly comparable to that for MTDT except for maximum queue length estimation. This is because the formation of queues at the location of each intersection follows a more complex pattern involving several intersection-level factors. The benefit of multi-task learning can be proved by comparing the maximum queue length estimation by MTDT-MOE which is designed to estimate MOEs exclusively and MTDT-Single. Supporting the performance made for queue length by the output of primary tasks, results in better estimation results.
As shown in MTDT results, the minimal marginal increase in error values for all variables remains minimal relative to the aggregation level indicating consistent reliability and robustness in the model’s performance across varying data resolutions. We also note the model exhibits enhanced performance in estimating travel time for lower quantiles of the data distribution while showing a slight improvement, particularly at the 90th quantile. This suggests that our designed digital twin is also particularly adept at capturing relatively rare or extreme patterns.
@@
TABLE 2: Performance of MTDT. The performance of our proposed digital twin is assessed across various aggregation resolutions of maximum queue length and at different quantiles of the travel time distribution. MTDT is compared to its variant (MTDT-Single and MTDT-MOE). This comparison is conducted in terms of NRMSE, RMSE, and MAE, measured in vehicle-per-bucket for Exit and Inflow waveforms, meter-per-bucket for maximum queue length, and vehicle-per-quantile for travel time distribution. Moreover, the level of confidence for estimating the relevant variable across various resolution levels is shown as 95% confidence intervals (CI).
For the multifaceted and fine-grained results, our designed multi-tasking simulation stands as unique to its counterparts, thus not directly comparable to existing baselines. However, we attempt to contrast our findings with those obtained by TQAM designed to make an accurate estimation of queue length estimation in a specific approach of a specific topology of target intersection. To achieve this, we partition our data into the same queue length buckets as reported for this model and compare error values in terms of (percentage) Mean Absolute Error (MAE) in Table 3. The partitioning encompasses two groups for each of the low (L), medium (M), and high (H) ranges of maximum queue length. These groups, ranging from L1 to H2, encompass 28%, 34%, 19%, 11%, 5%, and 3% of data samples, respectively. Our observations indicate that both MTDT and MTDT-Single exhibit comparable performance to TQAM. Specifically, MTDT outperforms TQAM in certain partitions of data where very short (0-40 meters) or long (above 200 meters) queue length tails are recorded, while MTDT-Single consistently performs equivalently or better than TQAM across nearly all partitions.
@@
TABLE 4: Evaluation at various green time allocations. The performance of the model is evaluated at different plit of data based on corridor-through green time percentage buckets.
@@ VI Related Work
In the realm of non-linear models, multi-kernel SVM [10] and weighted k-nearest neighbor [21] with MapReduce algorithm on Hadoop are introduced. When it comes to hybrid models, ensemble empirical mode decomposition with adaptive noise [16] and bidirectional RNN combined with road crossing vector coding [26] can be mentioned. Machine learning models in this line of research are deep learning frameworks based on the temporal convolutional network (TCN) [27], parallel computing of deep belief networks (DBNs) [25], deep characteristic learning using DBN with the multi-objective particle swarm optimization (PSO) algorithm [9]. [3] and [11] developed a deep hybrid neural network combined with CNN and LSTM using trajectory data for better accuracy and robustness of forecasting performance, while [23] combined GCN and attention mechanism for lane-wise and topology invariant traffic flow estimation.
@@ VII Conclusions and Future Work
In this paper, we demonstrate the effectiveness of multi-tasking as a learning paradigm for deep learning-based digital twins. We introduce a Multi-Task Deep Learning Digital Twin that utilizes multiple convolutions over graphs and time series. This approach can successfully perform a variety of related traffic simulation tasks for urban intersections with arbitrary topologies and characteristics. Our findings highlight the benefits of the multi-task learning approach in the design of our digital twin, not only for secondary tasks whose inputs rely on the output of primary tasks but also for primary tasks to generalize and enhance their performance through shared learned representations and joint optimization. Furthermore, we design two experiments to specifically illustrate how the model’s performance varies across different partitions of data when simulation scenarios are segmented based on signal timing plans or maximum recorded queue length. The results underscore the consistency and robustness of our approach across different experiments. Moreover, we discuss the lack of a baseline due to the uniqueness and comprehensiveness of our approach. However, we show it still outperforms in individual tasks compared to some previously introduced digital twins specialized for specific tasks.
@@
MTDT can serve as a valuable tool to assist traffic decision-makers in comprehending the potential impact of alterations to signal timing plans and proposed changes to road geometry. Additionally, it can be seamlessly integrated into signal timing optimization frameworks or expanded from isolated intersections to corridors and networks which will be a focal point of our future endeavors.